OpenStudy (anonymous):

While traveling to Europe, Phelan exchanged 250 US dollars for euros. He spent 150 euros on his trip. After returning to the United States he converts his money back to US dollars. How much of the original 250 US dollars does Phelan now have? 1 European euro = 1.3687 US dollars A. 44.70 US dollars B. 73.06 US dollars C. 136.87 US dollars D. 140.41 US dollars

OpenStudy (solomonzelman):

IS the course actually 1.36 ?

OpenStudy (dan815):

1 US dollar = 1/1.3687 of 1 European dollar

OpenStudy (solomonzelman):

If he changed dollars to euro in American Airport, then he got taxed like 30% additionaly

OpenStudy (anonymous):

umm???

OpenStudy (solomonzelman):

Courses can be different For example in America you can get 8 peso for 1$, where as in Argentina you can get 12 or close to 12 peso. (I've just been to Argentina :) )

OpenStudy (anonymous):

im on a timed quiz and i really need help

OpenStudy (dan815):

150 Euros Spent = 150/1.3678 US Dollar US DOLLA left with 250-150/1.3687=

OpenStudy (anonymous):

i got 140.4??

OpenStudy (anonymous):

oh! 140.41 then??

OpenStudy (dan815):

looks about right

OpenStudy (anonymous):

oh ok thanks :)